# MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations

### v1

@@ 7 Limitations
While MIRAGE provides a comprehensive evaluation framework for multimodal expert-level reasoning within agricultural consultations, several inherent limitations remain. First, the scope of the benchmark primarily reflects smallholder and backyard gardening scenarios, sourced largely from AskExtension’s database. Consequently, it does not fully encompass large-scale industrial agricultural practices or the more complex logistical, economic, and environmental issues prevalent in commercial farming operations. Additionally, agriculture is an extensive and multidisciplinary domain, covering areas such as soil science, agricultural engineering, economics, environmental sustainability, and policy implications. The current dataset captures only a subset of these dimensions. In the multi-turn scenario (MMMT), although designed to assess models’ ability to reason over dialogue context, the task does not yet simulate dynamic, truly interactive exchanges between models and users. The evaluations focus primarily on the models’ capacity to interpret and infer latent contextual information and to determine appropriate conversational actions based on partial observability. Consequently, MIRAGE does not fully measure interactive conversational reasoning, such as handling conversational repairs or real-time user feedback.
@@ Appendix A Data Source
The MIRAGE benchmark is constructed from a large-scale archive of real-world agricultural consultations obtained from Ask Extension [25], a U.S. national digital platform maintained by the Extension Foundation. Ask Extension is part of the broader Cooperative Extension System, a federally supported network of land-grant universities that delivers science-based, community-oriented education and services across the United States. The platform connects members of the public, such as farmers, gardeners, or homeowners, with university-affiliated experts who provide timely, research-backed responses to their questions.
@@
As seen in Figure 3, Our filtered AskExtension dialogues are strongly multimodal and reference-rich. Users typically include one to three images in their questions, about 35% of turns have a single image, 26% include two, and 39% include three. Experts in turn ground their advice in external sources: roughly 70% of answers cite a single URL, while the remaining 30% provide two. The total amount of content fetched from those links spans from about 2 000 up to 10 000 tokens per response, indicating that experts draw on substantial external context to support their guidance.
@@ Appendix B Related Works
Multimodal Large Language Models: Recent advances in multimodal large language models (LLMs) have markedly expanded vision–language reasoning capabilities. Proprietary models such as GPT-4 [8], Claude 3 Sonnet [10], and Google’s Gemini [55] demonstrate strong capabilities in unifying visual and textual modalities, achieving notable success across diverse multimodal benchmarks. Concurrently, open-source models—including Qwen-VL 2.5 [14], Gemma 3 [56], and InternVL-3 [73]—have narrowed the performance gap while remaining publicly accessible. Although these models excel on general-domain benchmarks, they underperform in agriculture: they lack fine-grained visual expertise, agronomic terminology, and the ability to reason about rare biological entities and management practices. MIRAGE is designed to expose these weaknesses by providing domain-specific, multimodal tasks that require expert-level diagnosis and advice.
@@
Multimodal Agricultural Benchmarks: Agricultural benchmarks [19, 26, 31] have progressed from controlled, image-only datasets to more realistic vision–language corpora. Early resources such as PlantVillage [36] and PlantDoc [52] emphasise leaf-level disease classification under laboratory and in-field conditions, respectively. The AgMMU [30] benchmark augments field images with farmer–expert dialogues, yet its evaluation is dominated by short, synthetic multiple-choice questions that limit open-ended reasoning. The recent Crop Disease Domain Multimodal (CDDM) [42] corpus contributes 137k crop–disease images paired with 1 M single-turn QA pairs, but it focuses narrowly on disease identification and management. In contrast, MIRAGE is derived from large-scale, real-world consultations covering more than 7000 biological entities; it offers both single-turn and multi-turn tasks, explicitly models clarify-or-respond decisions, and scores long-form answers along multiple dimensions of reasoning quality, thereby providing a far more ecologically valid and challenging test bed.
@@
Long-Form Question Answering Evaluation: Lexical-overlap metrics such as ROUGE [39], BLEU [46], and BERTScore [72] are ill-suited to evaluating open-ended, knowledge-intensive answers, as they correlate poorly with human judgements. Recent work therefore adopts LLM-as-Judge paradigms, exemplified by AlpacaEval [22], MT-Bench [13], and G-Eval [43], in which powerful LLMs are prompted to assess responses along axes such as factuality, coherence, and completeness. While these methods better align with human preferences, single-model evaluations remain vulnerable to bias and opacity. MIRAGE employs an interpretable ensemble of reasoning-focused LLM judges DeepSeek-R1-Distilled [34], Qwen 3-32B [5], and Phi-4-Reasoning [6] and by conducting three independent generation–evaluation passes per sample. This multi-model, multi-run protocol enhances robustness, enables variance analysis, and provides publicly inspectable rationales, yielding transparent and reproducible long-form QA evaluation.
@@
State Tracking in Dialogue Systems: State tracking has been a central component in task-oriented dialogue systems [16], where it plays a critical role in maintaining a representation of user intent and progress toward the user’s goal throughout the interaction. Dialogue state tracking (DST) typically involves predicting a belief state, a structured representation of slot-value pairs reflecting the current user intent at every turn. The widespread adoption of state tracking in dialogue systems has enabled robust multi-turn interactions, improved task success rates, and enhanced user satisfaction. Benchmarks like MultiWOZ [15, 24, 66] and Schema-Guided Dialogue (SGD) [47] have been instrumental in shaping the field by introducing multi-domain, open-schema challenges that pushed models to become more generalizable and scalable. These successes in dialogue suggest that structured, continuously updated representations of task progress can be highly beneficial in other interactive or sequential decision-making settings. Building on this insight, we introduce Goal State Tracking (GST), a mechanism that explicitly monitors whether all user-provided information is sufficient to generate a sound consultative recommendation; when gaps remain, GST triggers the agent to pose targeted clarifying questions, ensuring that advice is delivered only once the requisite context is complete.
@@ C.4 Biological Entity Synonymy Collection
To facilitate fair evaluation of biological entity identification in model outputs, we developed a comprehensive name collection pipeline that aggregates all valid references, both scientific and vernacular, for biological entities in our dataset. Rather than standardizing to a single canonical form, our objective was to compile an exhaustive synonymy for each entity, enabling robust matching regardless of the nomenclatural variant used. The system queries multiple authoritative taxonomic databases via their APIs in a hierarchical approach, beginning with the Global Biodiversity Information Facility (GBIF) [2] and extending to iNaturalist [58] Encyclopedia of Life (EOL) [1], Integrated Taxonomic Information System (ITIS) [3], Wikipedia [62], and NCBI Taxonomy [4]. Queries are enhanced with category-specific context (e.g., kingdom Plantae for botanical entries) to improve retrieval accuracy. For each entity, we preserve all retrieved scientific names (including accepted names, synonyms, and historical nomenclature) and vernacular names across languages, with metadata indicating the source authority. This comprehensive approach prevents unfair penalization during model evaluation when, for instance, a model correctly identifies an organism using its scientific name (Phytolacca americana) while the reference answer uses a common name ("pokeweed"), or vice versa. The resulting enriched dataset maintains the original hierarchical structure while appending all valid nomenclatural alternatives, thereby supporting more equitable assessment of biological entity recognition capabilities.
@@ D.1.1 Traditional LLM-as-Judge
Traditional metrics for evaluating long-form question answering have significant limitations. Research by Xu et al. [64] reveals that established approaches like ROUGE and BERTScore frequently diverge from human quality assessments, creating a fundamental evaluation challenge. A breakthrough solution has emerged in the form of language model-based evaluation frameworks. Cortes et al. [21] found that carefully designed prompting strategies with advanced models like GPT-4 can achieve remarkable alignment with human judgment, particularly when assessing the thoroughness of responses. This paradigm shift suggests that AI systems themselves may offer the most effective tools for evaluating complex natural language generation tasks.
LLMs have become widely adopted as evaluators in benchmarks like AlpacaEval [23], MT-Bench [13], and Chatbot Arena [18], where models such as GPT-4 conduct pairwise preference assessments based on helpfulness, factuality, and engagement. Frameworks like G-Eval [43] further enable fine-grained scoring by prompting models to assess specific dimensions using structured rubrics. These LLM-based evaluations have shown stronger alignment with human judgments than traditional metrics, especially on long-form and knowledge-intensive tasks.
@@ D.1.2 Leveraging an Interpretable Ensemble of Reasoning LLMs
To address the limitations of single-model, single-pass evaluation pipelines, we propose an interpretable and robust evaluation framework based on an ensemble of reasoning-capable LLMs: Deepseek-R1-Distilled [34], Qwen-3-32B [5], and Phi-4-Reasoning [6]. These models were selected not only for their demonstrated strength in multi-hop reasoning and long-form comprehension but also for their open accessibility, ensuring that our evaluation protocol is fully transparent and reproducible without dependence on proprietary APIs. Our evaluation protocol is also distinguished by its multi-run robustness. For each benchmark sample, we perform three independent inference runs of the candidate model to capture natural generation variability. Each of these outputs is then evaluated independently by the full ensemble of three judge models. This results in nine total evaluations per sample, allowing us to report aggregated metrics that reflect not only average performance but also stability and consistency across generations and evaluators. We further analyze cross-model agreement and judgment variance to ensure evaluation fidelity. This interpretable ensemble-based and multi-run evaluation represents a significant step forward in the use of LLMs as evaluators. It brings together the benefits of scale and automation while maintaining experimental rigor.
@@ D.2 Reliability and Robustness of Multi-Judge Evaluation
Fleiss’ Kappa: To measure categorical agreement across multiple judges, we binarize evaluation outcomes (e.g., correct vs. incorrect) and compute Fleiss’ [28], a widely used metric for evaluating agreement on nominal data among fixed raters. This quantifies how consistently the LLM judges classify outputs beyond what would be expected by chance.
Kendall’s W (Coefficient of Concordance: For tasks involving ordinal scoring or ranking of generations, we compute Kendall’s W [27], a non-parametric measure of rank correlation. This accounts for judges using different scoring scales by focusing on relative orderings. To accommodate tied ranks, we use the corrected-for-ties version of Kendall’s W. The coefficient ranges from 0 (no agreement) to 1 (perfect agreement), enabling us to quantify the degree of concordance in evaluative judgments.
@@
For this purpose, we compute the Intraclass Correlation Coefficient (ICC), which quantifies how consistently a single model scores the same item across multiple runs. ICC is particularly suited for this setting as it accounts for both within-subject and between-subject variability, providing a continuous-scale assessment of intra-rater consistency. Following established interpretative guidelines [20] we classify ICC values into bands (e.g., moderate, good, excellent) to report the strength of reliability for each judge.
@@
D.3 Diagnostic Parsimony
Diagnostic parsimony refers to the principled restraint in offering explanations, favoring the simplest account consistent with observed evidence while avoiding unnecessary speculation. [45] shows that humans have a robust bias toward simpler explanations, not just for cognitive ease but because simplicity aids understanding, memory, and decision-making. In agriculture, this principle is especially critical: farmers and gardeners often seek immediate, actionable advice under time-sensitive or resource-constrained conditions. Overly elaborate responses can obscure key insights, introduce confusion, or even lead to misapplied interventions. As in medicine and law, where expert communication must balance completeness with clarity, parsimony is a cornerstone of effective agronomic consultation. LLM/LVLM answers in such knowledge-intensive domains must therefore strive to be concise yet comprehensive, ensuring the user gets just enough information to act confidently. Studies show that human evaluators, particularly non-experts, tend to prefer shorter, to-the-point responses, even when some details are omitted [65].
However, many instruction-following LLMs/LVLMs often default to verbose outputs, including tangential or overly cautious elaborations. As illustrated in Figure 8, when responding to a user’s concern about bark defects on a tree, GPT-4.1 generates a detailed and structured diagnostic explanation covering wound physiology, tree tolerance, preventative care, and multi-step recommendations. Qwen 2 VL 72B, while more concise, still includes speculative possibilities (e.g., fungal infection, water stress) and a broad checklist of actions. In contrast, the human expert offers a brief but decisive interpretation, attributing the damage to non-serious activity and explicitly recommending no treatment. This behavior not only hinders user trust and satisfaction but also confounds traditional QA metrics like exact match (EM) and F1, which are poorly suited to judge informativeness under parsimony constraints. [9].

### v2

@@ 7 Limitations & Future Work
While MIRAGE provides a comprehensive evaluation framework for multimodal expert-level reasoning within agricultural consultations, several opportunities for future improvement remain. Currently, the benchmark primarily reflects smallholder and backyard gardening scenarios, largely derived from the AskExtension database. In future work, we aim to expand its scope to better represent large-scale industrial agricultural practices and the broader logistical, economic, and environmental challenges characteristic of commercial farming operations. Moreover, agriculture is inherently multidisciplinary, encompassing soil science, agricultural engineering, economics, environmental sustainability, and policy considerations. Future iterations of MIRAGE will extend coverage across these dimensions to better capture the full breadth of agricultural reasoning. In the multi-turn scenario (MMMT), our present design assesses models’ ability to reason over dialogue context, but it does not yet fully simulate dynamic, interactive exchanges between models and users. To address this, future versions will incorporate agentic capabilities that enable models to leverage time- and location-based context, as well as respond adaptively to evolving user input. Such extensions will allow MIRAGE to better evaluate interactive conversational reasoning, including conversational repairs, user feedback integration, and contextually grounded decision-making.
@@ 9 Acknowledgments
MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations
@@ Appendix A Data Source
The MIRAGE benchmark is constructed from a large-scale archive of real-world agricultural consultations obtained from Ask Extension [extension_ask_extension], a U.S. national digital platform maintained by the Extension Foundation. Ask Extension is part of the broader Cooperative Extension System, a federally supported network of land-grant universities that delivers science-based, community-oriented education and services across the United States. The platform connects members of the public, such as farmers, gardeners, or homeowners, with university-affiliated experts who provide timely, research-backed responses to their questions.
@@
As seen in Figure 4, Our filtered AskExtension dialogues are strongly multimodal and reference‐rich. Users typically include one to three images in their questions, about 35% of turns have a single image, 26% include two, and 39% include three. Experts in turn ground their advice in external sources: roughly 70% of answers cite a single URL, while the remaining 30% provide two. The total amount of content fetched from those links spans from about 2 000 up to 10 000 tokens per response, indicating that experts draw on substantial external context to support their guidance.
@@ Appendix B Related Works
Multimodal Large Language Models: Recent advances in multimodal large language models (LLMs) have markedly expanded vision–language reasoning capabilities. Proprietary models such as GPT-4 [gpt-4], Claude 3 Sonnet [claude-sonnet], and Google’s Gemini [team2023gemini] demonstrate strong capabilities in unifying visual and textual modalities, achieving notable success across diverse multimodal benchmarks. Concurrently, open‑source models—including Qwen‑VL 2.5 [bai2025qwen2], Gemma 3 [team2025gemma], and InternVL‑3 [zhu2025internvl3] have narrowed the performance gap while remaining publicly accessible. Although these models excel on general‑domain benchmarks, they underperform in agriculture: they lack fine‑grained visual expertise, agronomic terminology, and the ability to reason about rare biological entities and management practices. MIRAGE is designed to expose these weaknesses by providing domain‑specific, multimodal tasks that require expert‑level diagnosis and advice.
@@
Multimodal Agricultural Benchmarks: Agricultural benchmarks [agri_bench1, agri_bench2, agri_bench3] have progressed from controlled, image‑only datasets to more realistic vision–language corpora. Early resources such as PlantVillage [PlantVillage] and PlantDoc [singh2020plantdoc] emphasise leaf‑level disease classification under laboratory and in‑field conditions, respectively. The AgMMU [gauba2025agmmu] benchmark augments field images with farmer–expert dialogues, yet its evaluation is dominated by short, synthetic multiple‑choice questions that limit open‑ended reasoning. The recent Crop Disease Domain Multimodal (CDDM) [CDDM] corpus contributes 137k crop–disease images paired with 1 M single‑turn QA pairs, but it focuses narrowly on disease identification and management. In contrast, MIRAGE is derived from large‑scale, real‑world consultations covering more than 7000 biological entities; it offers both single‑turn and multi‑turn tasks, explicitly models clarify‑or‑respond decisions, and scores long‑form answers along multiple dimensions of reasoning quality, thereby providing a far more ecologically valid and challenging test bed.
@@
Long‑Form Question Answering Evaluation: Lexical‑overlap metrics such as ROUGE [rouge], BLEU [bleu], and BERTScore [zhang2019bertscore] are ill‑suited to evaluating open‑ended, knowledge‑intensive answers, as they correlate poorly with human judgements. Recent work therefore adopts LLM‑as‑Judge paradigms, exemplified by AlpacaEval [alpacaeval], MT‑Bench [bai2024mt], and G‑Eval [liu2023g], in which powerful LLMs are prompted to assess responses along axes such as factuality, coherence, and completeness. While these methods better align with human preferences, single‑model evaluations remain vulnerable to bias and opacity. MIRAGE employs an interpretable ensemble of reasoning‑focused LLM judges DeepSeek‑R1‑Distilled [guo2025deepseek], Qwen 3‑32B [githubQwen3Qwen3_Technical_ReportpdfMain], and Phi‑4‑Reasoning [abdin2025phi] and by conducting three independent generation–evaluation passes per sample. This multi‑model, multi‑run protocol enhances robustness, enables variance analysis, and provides publicly inspectable rationales, yielding transparent and reproducible long‑form QA evaluation.
@@
State Tracking in Dialogue Systems: State tracking has been a central component in task-oriented dialogue systems [chen2017survey, liu2024convbench], where it plays a critical role in maintaining a representation of user intent and progress toward the user’s goal throughout the interaction. Dialogue state tracking (DST) typically involves predicting a belief state, a structured representation of slot-value pairs reflecting the current user intent at every turn. The widespread adoption of state tracking in dialogue systems has enabled robust multi-turn interactions, improved task success rates, and enhanced user satisfaction. Benchmarks like MultiWOZ [budzianowski2018multiwoz, eric2019multiwoz, ye2021multiwoz] and Schema-Guided Dialogue (SGD) [rastogi2020towards] have been instrumental in shaping the field by introducing multi-domain, open-schema challenges that pushed models to become more generalizable and scalable. These successes in dialogue suggest that structured, continuously updated representations of task progress can be highly beneficial in other interactive or sequential decision-making settings. Building on this insight, we introduce Goal State Tracking (GST), a mechanism that explicitly monitors whether all user-provided information is sufficient to generate a sound consultative recommendation; when gaps remain, GST triggers the agent to pose targeted clarifying questions, ensuring that advice is delivered only once the requisite context is complete.
@@ C.4 Biological Entity Synonymy Collection
To facilitate fair evaluation of biological entity identification in model outputs, we developed a comprehensive name collection pipeline that aggregates all valid references, both scientific and vernacular, for biological entities in our dataset. Rather than standardizing to a single canonical form, our objective was to compile an exhaustive synonymy for each entity, enabling robust matching regardless of the nomenclatural variant used. The system queries multiple authoritative taxonomic databases via their APIs in a hierarchical approach, beginning with the Global Biodiversity Information Facility (GBIF) [GBIF] and extending to iNaturalist [van2018inaturalist] Encyclopedia of Life (EOL) [EOL], Integrated Taxonomic Information System (ITIS) [ITISgov], Wikipedia [Wikipedia], and NCBI Taxonomy [NIHTaxonomy]. Queries are enhanced with category-specific context (e.g., kingdom Plantae for botanical entries) to improve retrieval accuracy. For each entity, we preserve all retrieved scientific names (including accepted names, synonyms, and historical nomenclature) and vernacular names across languages, with metadata indicating the source authority. This comprehensive approach prevents unfair penalization during model evaluation when, for instance, a model correctly identifies an organism using its scientific name (Phytolacca americana) while the reference answer uses a common name ("pokeweed"), or vice versa. The resulting enriched dataset maintains the original hierarchical structure while appending all valid nomenclatural alternatives, thereby supporting more equitable assessment of biological entity recognition capabilities.
@@ D.1.1 Traditional LLM-as-Judge
Traditional metrics for evaluating long-form question answering have significant limitations. Research by Xu et al. [criticalxu2023critical] reveals that established approaches like ROUGE and BERTScore frequently diverge from human quality assessments, creating a fundamental evaluation challenge. A breakthrough solution has emerged in the form of language model-based evaluation frameworks. Cortes et al. [beyondcortes2024beyond] found that carefully designed prompting strategies with advanced models like GPT-4 can achieve remarkable alignment with human judgment, particularly when assessing the thoroughness of responses. This paradigm shift suggests that AI systems themselves may offer the most effective tools for evaluating complex natural language generation tasks.
LLMs have become widely adopted as evaluators in benchmarks like AlpacaEval [dubois2024length], MT-Bench [bai2024mt], and Chatbot Arena [chiang2024chatbot], where models such as GPT-4 conduct pairwise preference assessments based on helpfulness, factuality, and engagement. Frameworks like G-Eval [liu2023g] further enable fine-grained scoring by prompting models to assess specific dimensions using structured rubrics. These LLM-based evaluations have shown stronger alignment with human judgments than traditional metrics, especially on long-form and knowledge-intensive tasks.
@@ D.1.2 Leveraging an Interpretable Ensemble of Reasoning LLMs
To address the limitations of single-model, single-pass evaluation pipelines, we propose an interpretable and robust evaluation framework based on an ensemble of reasoning-capable LLMs: Deepseek-R1-Distilled [guo2025deepseek], Qwen-3-32B [githubQwen3Qwen3_Technical_ReportpdfMain], and Phi-4-Reasoning [abdin2025phi]. These models were selected not only for their demonstrated strength in multi-hop reasoning and long-form comprehension but also for their open accessibility, ensuring that our evaluation protocol is fully transparent and reproducible without dependence on proprietary APIs. Our evaluation protocol is also distinguished by its multi-run robustness. For each benchmark sample, we perform three independent inference runs of the candidate model to capture natural generation variability. Each of these outputs is then evaluated independently by the full ensemble of three judge models. This results in nine total evaluations per sample, allowing us to report aggregated metrics that reflect not only average performance but also stability and consistency across generations and evaluators. We further analyze cross-model agreement and judgment variance to ensure evaluation fidelity. This interpretable ensemble-based and multi-run evaluation represents a significant step forward in the use of LLMs as evaluators. It brings together the benefits of scale and automation while maintaining experimental rigor.
@@ D.2 Reliability and Robustness of Multi-Judge Evaluation
Fleiss’ Kappa: To measure categorical agreement across multiple judges, we binarize evaluation outcomes (e.g., correct vs. incorrect) and compute Fleiss’ [fleiss1971measuring], a widely used metric for evaluating agreement on nominal data among fixed raters. This quantifies how consistently the LLM judges classify outputs beyond what would be expected by chance.
Kendall’s W (Coefficient of Concordance: For tasks involving ordinal scoring or ranking of generations, we compute Kendall’s W [field2005k], a non-parametric measure of rank correlation. This accounts for judges using different scoring scales by focusing on relative orderings. To accommodate tied ranks, we use the corrected-for-ties version of Kendall’s W. The coefficient ranges from 0 (no agreement) to 1 (perfect agreement), enabling us to quantify the degree of concordance in evaluative judgments.
@@
For this purpose, we compute the Intraclass Correlation Coefficient (ICC), which quantifies how consistently a single model scores the same item across multiple runs. ICC is particularly suited for this setting as it accounts for both within-subject and between-subject variability, providing a continuous-scale assessment of intra-rater consistency. Following established interpretative guidelines [cicchetti1994guidelines] we classify ICC values into bands (e.g., moderate, good, excellent) to report the strength of reliability for each judge.
@@
D.3 Management (MG) Evaluation Criteria and W-Sum Score
To evaluate model performance on the Management (MG) subset, we adopt a four-dimensional evaluation framework capturing complementary aspects of answer quality: Accuracy, Relevance, Completeness, and Parsimony. These dimensions were chosen based on findings from the long-form QA literature and insights from agricultural experts.
Accuracy measures the factual correctness of the recommendation or diagnosis.
Relevance assesses whether the response remains on-topic and directly addresses the user’s query.
Completeness evaluates whether essential details (e.g., causal factors, conditions, or next-step actions) are covered.
Parsimony rewards concise, evidence-based explanations that avoid unnecessary speculation or verbosity.
The first three dimensions align with multi-axis evaluation principles in long-form QA studies [xu2023critical, beyondcortes2024beyond], which demonstrate that completeness and relevance correlate more strongly with human judgments than surface-level metrics such as ROUGE or BERTScore. The fourth dimension, Parsimony, is inspired by the cognitive and philosophical principle of simplicity (Occam’s razor), emphasizing clear and minimal explanations in expert communication. We provide a detailed discussion and illustrative examples of diagnostic parsimony in Appendix D.4.
To enable model comparison with a single scalar, we compute a Weighted-Sum (W-Sum) score that aggregates these four dimensions. The adopted weight ratio of 2:1:1:1 reflects expert consensus that factual Accuracy should carry the greatest influence, since factual errors can lead to misleading or harmful recommendations. While we also report per-dimension results for transparency, the W-Sum provides a concise, interpretable measure for aggregate ranking.
D.4 Diagnostic Parsimony
Diagnostic parsimony refers to the principled restraint in offering explanations, favoring the simplest account consistent with observed evidence while avoiding unnecessary speculation. [pacer2017ockham] shows that humans have a robust bias toward simpler explanations, not just for cognitive ease but because simplicity aids understanding, memory, and decision-making. In agriculture, this principle is especially critical: farmers and gardeners often seek immediate, actionable advice under time-sensitive or resource-constrained conditions. Overly elaborate responses can obscure key insights, introduce confusion, or even lead to misapplied interventions. As in medicine and law, where expert communication must balance completeness with clarity, parsimony is a cornerstone of effective agronomic consultation. LLM/LVLM answers in such knowledge-intensive domains must therefore strive to be concise yet comprehensive, ensuring the user gets just enough information to act confidently. Studies show that human evaluators, particularly non-experts, tend to prefer shorter, to-the-point responses, even when some details are omitted [xu2023critical].
However, many instruction-following LLMs/LVLMs often default to verbose outputs, including tangential or overly cautious elaborations. As illustrated in Figure 8, when responding to a user’s concern about bark defects on a tree, GPT-4.1 generates a detailed and structured diagnostic explanation covering wound physiology, tree tolerance, preventative care, and multi-step recommendations. Qwen 2 VL 72B, while more concise, still includes speculative possibilities (e.g., fungal infection, water stress) and a broad checklist of actions. In contrast, the human expert offers a brief but decisive interpretation, attributing the damage to non-serious activity and explicitly recommending no treatment. This behavior not only hinders user trust and satisfaction but also confounds traditional QA metrics like exact match (EM) and F1, which are poorly suited to judge informativeness under parsimony constraints. [adlakha2024evaluating].